Consider two neighboring island countries called Contente and Euphoria. They each have 4 million labor hours available per moContentes opportunity cost of producing 1 bushel of corn is 1/2 pair of jeans, and Euphorias opportunity cost of producingCalculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relativeUnsure if the options selected are correct or not.
Answer #1

Blanks-

1) 16/8 = 2 pair of jeans

2) 20/5 =4 pair of jeans

3) Contente

4) Euphoria

5) 8*4 =32

6) 20*4 =80

7) 32-23 = 9

8) 80-68 =12

Contente

Euphoria

Corn

Jeans

Corn

Jeans

Without Trade

Production

8

48 15 20

Consumption

8

48 15 20

With Trade

Production

32

0

0

80

Trade action

Export 18

Import 54

Import 18

Export 54

Consumption

14

54

18

26

Gains from Trade

Increase in Consumption

6

6 3 6